A particle with the same charge as an electron experiences a force of 5.2 × 10-12N and traveling at a speed of 4.21 × 107m/s at right angles to a magnetic field. a. How strong is the magnetic field? b. What does its acceleration if its mass is 1.88 × 10-28 kg?

1
Expert's answer
2022-04-11T12:07:20-0400

a.

"B=\frac{F}{qv}=\frac{5.2*10^{-12}}{1.6*10^{-19}*4.21*10^7}=0.077\:\rm T"

b.

"a=\frac{F}{m}=\frac{5.2*10^{-12}}{1.88*10^{-28}}=2.77*10^{16}\:\rm m/s^2"